OpenStudy (anonymous):

a line with a slope of -3 crosses the x-axis at (8,0). At what point does it cross the y-axis?

OpenStudy (tkhunny):

General Slope Interecpt y = mx + b We are given the slope y = -3x + b We are given a point. 0 = -3(8) + b One thing at a time. substitute what you know and go looking for whatever is left.

OpenStudy (anonymous):

slope -3 means the line fall(- sign), and ration of rise/run is 3 to 1. imagine the graph in mind. passed (8,0) means run 8 units on X axis. if slope ratio 3/1 = n/8 means 1 become 8, and 3 become N. so N=3x8=24.